# Bayesian Density-Density Regression with Application to Cell-Cell Communications

### v1

@@ Appendix A MCMC
Appendix B Graph Discovery
In Figures 7-10 we show fitted means under Bayesian MLR and fitted densities under Bayesian DDR for all ordered pairs of cell types. As before, we use PCA for the response distribution of each donor to obtain a 2-dimensional projection. We observe that the predictions align with the reported RPEs shown in Figure 8 in the main text. Notably, Bayesian DDR provides more precise fits compared to Bayesian MLR, especially in cases where the response distributions are multimodal or exhibit high uncertainty.
@@ Appendix B Graph Discovery
Figure 8 shows boxplots of RPE for all ordered pairs of cell types. As before, we use the intercept model as the reference $f_{0}$ for computing RPEs. Figures 11-22 show boxplots for the linear regression coefficients of all ordered pairs of cell types. Such inference allows us to identify which ligands are significant in predicting a receptor. This is done by inspecting the posterior distribution for the corresponding coefficient $A_{eij}$. Values far from zero indicate significant receptor/ligand pairs. Consider, for example, the regression of B cell receptor expression on T cell ligands. Both ligands, CD40L and IFN-$\gamma$ are significant in predicting receptors such as IGHM, IGHD, IGHG, and IL-6R. Specifically, the ligand IFN-$\gamma$ appears critical to predict gene expression for the receptor IGHA, while CD40L is required to predict receptors CD40, ICOSL, IL-21R, BAFFR, CXCR5, and CCR7. Similar conclusions can be drawn for other ligand/receptor relations. For a more formal assessment one could use pseudo inclusion probabilities like (16) to quantify the evidence for a ligand being significant in predicting a receptor.
@@
Appendix C Semi-Simulation

### v2

@@ Appendix B MCMC
Appendix C Graph Discovery
In Figures 8-11 we show fitted means under Bayesian MLR and fitted densities under Bayesian DDR for all ordered pairs of cell types. As before, we use PCA for the response distribution of each donor to obtain a 2-dimensional projection. We observe that the predictions align with the reported RPEs shown in Figure 9 in the main text. Notably, Bayesian DDR provides more precise fits compared to Bayesian MLR, especially in cases where the response distributions are multimodal or exhibit high uncertainty.
@@ Appendix C Graph Discovery
Figure 9 shows boxplots of RPE for all ordered pairs of cell types. As before, we use the intercept model as the reference $f_{0}$ for computing RPEs. Figures 12-23 show boxplots for the linear regression coefficients of all ordered pairs of cell types. Such inference allows us to identify which ligands are significant in predicting a receptor. This is done by inspecting the posterior distribution for the corresponding coefficient $A_{eij}$. Values far from zero indicate significant receptor/ligand pairs. Consider, for example, the regression of B cell receptor expression on T cell ligands. Both ligands, CD40L and IFN-$\gamma$ are significant in predicting receptors such as IGHM, IGHD, IGHG, and IL-6R. Specifically, the ligand IFN-$\gamma$ appears critical to predict gene expression for the receptor IGHA, while CD40L is required to predict receptors CD40, ICOSL, IL-21R, BAFFR, CXCR5, and CCR7. Similar conclusions can be drawn for other ligand/receptor relations. For a more formal assessment one could use pseudo inclusion probabilities like (21) to quantify the evidence for a ligand being significant in predicting a receptor.
@@
Appendix D Semi-Simulation